# Token Erasure as a Footprint of Implicit Vocabulary Items in LLMs

### v1

@@ 2 Background
Figure 2: Test accuracy on CounterFact subject last tokens versus other tokens in the dataset for probes trained on Llama-2-7b hidden states ($n=5063$). $i$ represents the position being predicted (e.g., $i=-1$ is previous token prediction; $i=1$ is next-token prediction). We observe an “erasure” effect in last subject tokens that is not present for other types of tokens: these last subject tokens consistently “forget” about preceding tokens and themselves. Appendix A shows Llama-3-8b results and in-distribution performance on Pile tokens.
@@ 3.1 Method
If last token positions are so important (Section 2), then what do these representations encode? Perhaps the last hidden state directly stores information about other subject tokens (e.g., _Wars might contain some encoding for _Star in its hidden state). To test this hypothesis, we investigate hidden states for both Llama-2-7b and Llama-3-8b, as they have significantly different token vocabulary sizes (32k and 128k tokens, respectively). We train linear probes $p_{i}^{(\ell)}$ to take a hidden state $h_{t}^{(\ell)}$ at layer $\ell$ and token position $t$ and predict the value of a nearby token $t+i$. (e.g., a probe trained to predict the previous token for layer 5 hidden states would be denoted by $p_{-1}^{(5)}$).
@@ 3.2 CounterFact Subjects
We also observe the same phenomenon when testing on named entities identified by spaCy in Wikipedia articles (Appendix A, Figure 12), suggesting that this effect is not an artifact of the short templates found in the CounterFact dataset. Additionally, we consider whether this result is due to imbalances in probe training data, but this seems not to be the case either (Appendix B).
@@ 3.3 Multi-Token Words
Intuitively, the process of converting a multi-token sequence like [_n, ort, he, astern] into a meaningful representation of the word “northeastern” resembles the process of converting [_E, iff, el, Tower] into “Eiffel Tower.” We hypothesize that models treat multi-token words in the same way that they treat multi-token entities, and test our probes from Section 3.1 on multi-token words. After sampling 500 articles ($\sim$256k tokens) from the 20220301.en split of the Wikipedia dump (Foundation, 2022), we split by white-space to naively identify word boundaries. As predicted, we see the same “erasing” pattern for multi-token words that we do for multi-token entities (Appendix A, Figure 11). This suggests that they may be processed in a similar manner in early layers.
@@ 4.1 An Erasure Score
We first define an erasure score $\psi_{p,q}$, which is a heuristic designed around the intuition that the last token representation of a lexical item should exhibit a strong “erasing” effect for probe predictions from layer 1 to layer $L$ [^3] For both Llama-2-7b and Llama-3-8b we set $L=9$. The score quantifies “erasing” behavior at token positions within a given sequence at indices $p$ through $q$, and penalizes any erasure of tokens outside of these boundaries. Equation 1 defines the score $\psi_{p,q}$ for a sequence $s_{p,q}$ of length $n=q-p+1$ as:
@@
where $\delta(t,i)$ denotes the change in probability of the predicted token $t+i$ from layer 1 to layer $L$, based on probes $p_{i}^{(\ell)}$ from Section 3.1.
@@
If $t+i$ lies outside the boundaries of $s$, we decrease $\psi_{p,q}$. Otherwise, a large drop between layers $\delta(t,i)$ increases the value of $\psi_{p,q}$.
@@
We provide further explanation of the intuition behind this approach in Appendix C.
@@ 4.2 Segmenting Documents
We develop an algorithm built around our erasure score $\psi$ that breaks any given document $d\in\mathcal{D}$ into high-scoring, non-overlapping segments covering all of $d$ (Algorithm 1, Appendix D). Figure 1 shows the top-scoring sequences $s_{p,q}$ calculated in this manner from a Wikipedia excerpt about Thelonious Monk, where unigram scores are excluded for clarity. Not all multi-token words are scored highly via our approach, but the highest-scoring sequences are plausible lexical items that are non-compositional in nature (“dram.atic”, “sil.ences”, “tw.ists”). We share examples of complete segmentations in Appendix D.
4.3 Model Vocabularies
MTW MTE
llama data prec. recall prec. recall
2-7b wiki 0.306 0.016 0.143 0.016
pile 0.296 0.017 0.080 0.018
3-8b wiki 0.044 0.001 0.010 0.000
pile 0.023 0.001 0.012 0.001
Finally, we propose a method to “read out” the implicit vocabulary of a model $\mathcal{M}$ given a dataset $\mathcal{D}$. For each document $d\in\mathcal{D}$, we segment $d$ using Algorithm 1. We then average scores $\psi$ for every multi-token sequence that appears more than once across all documents. As this process is very data-dependent, we compare the top 50 results for Pile and Wikipedia text in Appendix E.
With this approach, we are able to recover $\sim$1800 sequences for Llama-2-7b and $\sim$900 for Llama-3-8b using the same five hundred Wikipedia articles. Although recall is quite low (Table 1), we find that 44.9% of sequences recovered for Llama-2-7b on Wikipedia text are either multi-token words or multi-token entities (29.68% for Pile text). For Llama-3-8b, only 5% and 3% of sequences are MTWs or MTEs. However, looking at examples of Llama-3-8b sequences in Appendix E, we can observe other interesting cases, like multi-token expressions (“gold medalists,” “by per capita income,” “thank you for your understanding”) and LaTeX commands (as similarly observed by Elhage et al. (2022)). Because Llama-3-8b’s token vocabulary is four times larger than Llama-2-7b’s, its implicit vocabulary also seems to consist of more multi-word expressions and chunks of code rather than multi-token words (Appendix E, Table 6).
5 Conclusion
Limitations
Ethics Statement
Acknowledgments
We thank David Smith, Bilal Chughtai, Chantal Shaib, Atticus Geiger, and Adrian Chang for helpful discussion and feedback throughout the course of this project. This work was supported in part by Open Philanthropy, and by the National Science Foundation (NSF) grant IIS-1901117.
Experiments were implemented using the nnsight library; many were run on the Center for AI Safety Compute Cluster. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the sponsors.
Appendix A Linear Probing on Hidden States
A.1 Llama-3-8b Results
CounterFact Accuracy
We share results analogous to Figure 2 but for Llama-3-8b, which shows a similar “erasure” pattern (Figure 8). Probes are tested only on prompts that Llama-3-8b answers correctly.
Multi-Token Word Accuracy
Figure 9 shows results for probes tested on the last token positions of multi-token words from Wikipedia (where “words” are determined by whitespace separation).
Multi-Token Entity Accuracy
Figure 10 shows results for probes tested on the last token positions of multi-token entities identified by spaCy, using the same dataset as A.1. We use spaCy’s named entity recognition pipeline to identify named entities. Because digits 0-9 are added to Llama-2-7b’s vocabulary, we filter out all classes relating to numbers (PERCENT, DATE, CARDINAL, TIME, ORDINAL, MONEY, QUANTITY), with the thought that these sequences may be treated differently at the detokenization stage.
A.2 Llama-2-7b Results
Multi-Token Word Accuracy
Figure 11 shows results from Llama-2-7b probes tested on multi-token words from Wikipedia as described in Section 3.3.
Multi-Token Entity Accuracy
A.3 Llama-2-7b Extras
Pile Accuracy
Comparison of Token Positions
Comparison of Subject Lengths
Appendix B Accounting for Possible Training Imbalance
Appendix C Intuition for $\psi$
C.1 Explanation of Equation 1
Our first assumption when designing $\psi$ is that if an LLM is processing a sequence of tokens corresponding to a lexical item (e.g., [Cal, g, ary]), the last token ary in that sequence should “forget itself” between layers 1 and $L=9$, according to the pattern we observe in Section 3. We measure this drop using probability scores from probe outputs for a single example. In plain English, the first term $\delta(q,0)$ in Equation 1 represents how much $P(t=\texttt{ary})$ drops between layer 1 and layer $L$ when probing the hidden representation for ary.
In the double summation term, we take into account probe predictions for tokens one ($i=-1$) or two ($i=-2$) positions before the current token. If there is a drop in probability between layer 1 and $L$ for these positions, we take this as evidence of token clumping.
However, we also observe (from Figure 13 and manual inspection) that probe predictions for token positions that lie outside the boundaries of a presumed lexical item are maintained across layers, or even promoted. This is clear from the $i=-1$ case in the leftmost plot for Figure 13, and was also a consistent pattern when we examined probe behavior over a number of example documents. Given this fact, we include the indicator function $\mathbbm{1}_{\text{within}}(t,i)$ to decrease $\psi$ in cases where erasure is happening outside the bounds of the given sequence.
C.2 Choice of $L$
We choose $L=9$ based on probe “erasure” behavior for Llama-2-7b and Llama-3-8b, particularly Figure 2. We also present a short ablation experiment for $L\in\{5,9,13,17,21\}$ with results in Table 2, which shows that other values of $L$ after “the drop” are roughly equivalent to $L=9$.
MTW MTE
$L$ prec. recall prec. recall
5 0.307 0.002 0.143 0.002
9 0.306 0.016 0.143 0.016
13 0.328 0.003 0.169 0.003
17 0.330 0.003 0.180 0.003
21 0.319 0.003 0.172 0.003
Appendix D Document Segmentation
@@
We show full document segmentations using Algorithm 1 for short excerpts from the same Wikipedia article in Figure 4 and Figure 5. Figure 6 and Figure 7 show more segmentations for a Pile document.
@@ Appendix E Model Vocabularies
Tables 3 through 6 show the top 50 highest-scoring multi-token sequences for Llama-2-7b and Llama-3-8b across either five hundred Wikipedia articles or five hundred Pile samples. Entries were filtered to show only sequences that appear more than once.
@@
Figure 11: Test accuracy of probes on last tokens of Wikipedia multi-token words for Llama-2-7b ($n=80606$, right). Test accuracy on all other tokens shown on the left. Similarly to Figure 2, we see an erasing effect that is not present for other types of tokens.
@@
1992 births 7 2 0.573139
19th-century 7 3 0.568861
dehydrogen 5 2 0.553029

### v2

@@ 2 Background
Figure 2: Top-1 test accuracy on CounterFact subject last tokens versus other tokens in the dataset for probes trained on Llama-2-7b hidden states ($n=5063$). $i$ represents the position being predicted (e.g., $i=-1$ is previous token prediction; $i=1$ is next-token prediction). We observe an “erasure” effect in last subject tokens that is not present for other types of tokens: these last subject tokens consistently “forget” about preceding tokens and themselves. Appendix A shows Llama-3-8b results and in-distribution performance on Pile tokens.
@@ 3.1 Method
If last token positions are so important (Section 2), then what do these representations encode? Perhaps the last hidden state directly stores information about other subject tokens (e.g., _Wars might contain some encoding for _Star in its hidden state). To test this hypothesis, we investigate hidden states for both Llama-2-7b and Llama-3-8b, as they have significantly different token vocabulary sizes $|\mathcal{V}|$ (32k and 128k tokens, respectively).
Let $d$ denote the hidden dimension of the model. We train linear probes $p_{i}^{(\ell)}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{|\mathcal{V}|}$ to take a hidden state $h_{t}^{(\ell)}\in\mathbb{R}^{d}$ at layer $\ell$ and token position $t$ and predict the value of a nearby token $t+i$. For example, a probe trained to predict the previous token for hidden states at layer 5 would be denoted by $p_{-1}^{(5)}$.
@@ 3.2 CounterFact Subjects
We also observe the same phenomenon when testing on named entities identified by spaCy in Wikipedia articles (Appendix A, Figure 12), suggesting that this effect is not an artifact of the short templates found in the CounterFact dataset. It also does not seem to be a result of any imbalances in probe training data (Appendix B).
@@ 3.3 Multi-Token Words
Intuitively, the process of converting a multi-token sequence like [_n, ort, he, astern] into a meaningful representation of the word “northeastern” resembles the process of converting [_E, iff, el, Tower] into “Eiffel Tower.” We hypothesize that models treat multi-token words in the same way that they treat multi-token entities, and test our probes from Section 3.1 on multi-token words. After sampling 500 articles ($\sim$256k tokens) from the 20220301.en split of the Wikipedia dump (Foundation, 2022), we split by white-space to naively identify word boundaries. As predicted, we see the same “erasing” pattern for multi-token words that we do for multi-token entities (Figure 3). This suggests that they may be processed in a similar manner in early layers. Appendix A shows similar results for Llama-3-8b.
Figure 3: Top-1 test accuracy of probes on last tokens of Wikipedia multi-token words for Llama-2-7b ($n=80606$). Accuracy on all other tokens shown on the left. We see an erasing effect for multi-token words, similar to the effect seen for CounterFact subjects in Figure 2.
@@ 4.1 An Erasure Score
Given some arbitrary sequence of tokens from indices $p$ through $q$, we want to design an erasure score that captures intuitions from Section 3. This score should be higher for sequences exhibiting token erasure (which we hypothesize to be lexical items like [_Cal, g, ary]), and lower for other types of token sequences (e.g., [_go _to, _Cal, g]). We design a metric $\psi_{p,q}$ that uses probe outputs from Section 3 to look for erasure effects betwen layer 1 and layer $L$ [^3] For both Llama-2-7b and Llama-3-8b we set $L=9$.
Concretely, Equation 1 defines the score $\psi_{p,q}$ for a sequence $s_{p,q}$ of length $n=q-p+1$ as:
@@
where $\delta(t,i)$ denotes the change in probability of the predicted token $t+i$ from layer 1 to layer $L$, based on probes $p_{i}^{(\ell)}$ from Section 3.1. We take the softmax of the probe outputs to obtain the probability of a specific token $t+i$ in Equation 2.
@@
Finally, if $t+i$ lies outside the boundaries of $s$, we want the score to decrease. If it is within the boundaries of $s$, we want a large drop between layers $\delta(t,i)$ to increase the value of $\psi_{p,q}$.
@@
In summary, for every token position $p\leq t\leq q$ and prediction offset $i\in\{-2,-1\}$, we measure the drop in the predicted probability of the correct token $t+i$ between layer 1 and layer $L$. The more that the probability of the correct answer decreases in early layers, the higher we score that sequence. However, if this “forgetting” occurs for tokens outside of the boundaries of $s$, we subtract that value from the overall score, effectively penalizing the sequence. This intuition comes from close inspection of probe behavior—for example, Figure 13 shows that there is no “forgetting” effect for $i=-1$ when probing the first token of CounterFact subjects. With this approach, we can also account for cases where $s$ is a subsequence of a larger lexical item: if the token g shows a forgetting effect for _Cal in [_Cal, g, ary], then the sequence [g, ary] would be penalized. Finally, $\delta(q,0)$ additionally rewards sequences in which the last token “forgets itself,” as seen in Figures 2 and 3. We then normalize by the total number of $\delta$ values considered, in order to account for differing sequence lengths.
@@ 4.2 Segmenting Documents
We develop an algorithm built around our erasure score $\psi$ that breaks any given document $d\in\mathcal{D}$ into high-scoring, non-overlapping segments covering all of $d$ (Algorithm 1). Figure 1 shows the top-scoring sequences $s_{p,q}$ calculated in this manner from a Wikipedia excerpt about Thelonious Monk, where unigram scores are excluded for clarity. Not all multi-token words are scored highly via our approach, but the highest-scoring sequences are plausible lexical items that are non-compositional in nature (“dram.atic”, “sil.ences”, “tw.ists”). We share examples of more documents with complete segmentations in Appendix D.
@@
4.3 Model Vocabularies
Token Sequence $n$ ct $\psi$
lower case 3 2 0.736012
storm 2 4 0.716379
excursion 4 2 0.713134
====… (72 ‘equals’ signs) 8 2 0.712982
Mom 3 2 0.706778
acre 3 2 0.629213
Subject 3 2 0.607172
ninth 3 2 0.606669
processing elements 3 2 0.599549
CVC 3 2 0.596735
Table 1: Top ten highest-scoring sequences for Llama-2-7b using a Pile subsample (1658 sequences recovered total). $n$ is the number of tokens in the sequence, and ‘ct’ represents occurrences of this segment. $\psi$ is averaged over all occurrences.
Finally, we propose a method to “read out” the implicit vocabulary of a model $\mathcal{M}$ given a dataset $\mathcal{D}$. For each document $d\in\mathcal{D}$, we segment $d$ using Algorithm 1. We then average scores $\psi$ for every multi-token sequence that appears more than once in $\mathcal{D}$. As this process is very data-dependent, we show results for both Pile and Wikipedia text. The top 50 sequences for each dataset and model are provided in Appendix E.
With this approach, we are able to recover $\sim$1800 sequences for Llama-2-7b and $\sim$900 for Llama-3-8b using the same five hundred Wikipedia articles. Although recall is quite low (Table 2), we find that 44.9% of sequences recovered for Llama-2-7b on Wikipedia text are either multi-token words or multi-token entities (29.68% for Pile text). For Llama-3-8b, only 5% and 3% of retrieved sequences are multi-token words or entities. However, looking at examples of Llama-3-8b sequences in Appendix E, we can observe other interesting cases, like multi-token expressions (“gold medalists,” “by per capita income,” “thank you for your understanding”) and LaTeX commands (as similarly observed by Elhage et al. (2022)). Because Llama-3-8b’s token vocabulary is four times larger than Llama-2-7b’s, its implicit vocabulary also seems to consist of larger multi-word expressions and chunks of code rather than multi-token words (Appendix E, Table 7).
5 Conclusion
MTW MTE
llama data prec. recall prec. recall
2-7b wiki 0.306 0.016 0.143 0.016
pile 0.296 0.017 0.080 0.018
3-8b wiki 0.044 0.001 0.010 0.000
pile 0.023 0.001 0.012 0.001
Limitations
Ethics Statement
Acknowledgments
We thank Koyena Pal, David Smith, Bilal Chughtai, Chantal Shaib, Atticus Geiger, and Adrian Chang for helpful discussion and feedback throughout the course of this project. This work was supported in part by Open Philanthropy, and by the National Science Foundation (NSF) grant IIS-1901117.
Experiments were implemented using the nnsight library (Fiotto-Kaufman et al., 2024). Many were run on the Center for AI Safety Compute Cluster. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the sponsors.
Appendix A Additional Probing Results
A.1 Llama-3-8b Results
CounterFact Accuracy
We share results analogous to Figure 2 for Llama-3-8b, which shows a similar “erasure” pattern (Figure 9). Probes are tested only on prompts that Llama-3-8b answers correctly.
Multi-Token Word Accuracy
Figure 10 shows results for Llama-3-8b probes tested on the last token positions of multi-token words from Wikipedia (where “words” are determined by whitespace separation).
Multi-Token Entity Accuracy
Figure 11 shows results for probes tested on the last token positions of multi-token entities identified by spaCy, using the same dataset that we do for multi-token words. We use spaCy’s named entity recognition pipeline to identify named entities. Because digits 0-9 are added to Llama-2-7b’s vocabulary, we filter out all classes relating to numbers (PERCENT, DATE, CARDINAL, TIME, ORDINAL, MONEY, QUANTITY), with the thought that these sequences may be treated differently at the detokenization stage.
A.2 Llama-2-7b Results
Multi-Token Entity Accuracy
Pile Accuracy
Comparison of Token Positions
Comparison of Subject Lengths
Appendix B Accounting for Possible Training Imbalance
Appendix C Choice of $L$
We choose $L=9$ based on probe behavior for Llama-2-7b and Llama-3-8b, particularly in Figures 2 and 3. Table 3 shows an additional ablation experiment for $L\in\{5,9,13,17,21\}$.
MTW MTE
$L$ prec. recall prec. recall
5 0.307 0.002 0.143 0.002
9 0.306 0.016 0.143 0.016
13 0.328 0.003 0.169 0.003
17 0.330 0.003 0.180 0.003
21 0.319 0.003 0.172 0.003
Appendix D Document Segmentation
We provide full document segmentations using Algorithm 1 for a short excerpt from a Wikipedia article in Figures 5 and 6. Figures 7 and 8 show segmentations for a Pile document.
@@ Appendix E Model Vocabularies
Tables 4 through 7 show the top 50 highest-scoring multi-token sequences for Llama-2-7b and Llama-3-8b across either five hundred Wikipedia articles or five hundred Pile samples. Entries were filtered to show only sequences that appear more than once.
@@
1992 births 7 2 0.573
19th-century 7 3 0.569
dehydrogen 5 2 0.553